On the effective cone of the moduli space of pointed rational curves

Brendan Hassett and Yuri Tschinkel

Abstract. We compute the effective cone of the moduli space of stable curves of genus zero with six marked points.

1. Introduction

For a smooth projective variety, Kleiman’s criterion for ample divisors states that the closed ample cone (i.e., the nef cone) is dual to the closed cone of effective curves. Since the work of Mori, it has been clear that extremal rays of the cone of effective curves play a special role in birational geometry. These correspond to certain distinguished supporting hyperplanes of the nef cone which are negative with respect to the canonical class. Contractions of extremal rays are the fundamental operations of the minimal model program.

Fujita [F] has initiated a dual theory, with the (closed) cone of effective divisors playing the central role. It is natural then to consider the dual cone and its generators. Those which are negative with respect to the canonical class are called coextremal rays, and have been studied by Batyrev [Ba]. They are expected to play a fundamental role in Fujita’s program of classifying fiber-space structures on polarized varieties.

There are relatively few varieties for which the extremal and coextremal rays are fully understood. Recently, moduli spaces of pointed rational curves \overline{M}_{0,n} have attracted considerable attention, especially in connection with mathematical physics and enumerative geometry. Keel and McKernan first considered the ‘Fulton conjecture’: The cone of effective curves of \overline{M}_{0,n} is generated by one-dimensional boundary strata. This is proved for n ≤ 7 [KeMc]. The analogous statement for divisors, namely, that the effective cone of \overline{M}_{0,n} is generated by boundary divisors, is known to be false ([Ke] and [Ve]). The basic idea is to consider the map

r : \overline{M}_{0,2g} ↪ \overline{M}_{g}, n = 2g,

identifying pairs (i_{1}i_{2}), (i_{3}i_{4}), …, (i_{2g−1}i_{2g}) of marked points to nodes. There exist effective divisors in \overline{M}_{g} restricting to effective divisors not spanned by boundary divisors (see Remark 4.2). However, it is true that for each n the cones of 𝔖_{n}- invariant effective divisors are generated by boundary divisors [KeMc].

1991 Mathematics Subject Classification. Primary 14E30; Secondary 14D22.

Key words and phrases. Effective cones, moduli spaces.

The first author was partially supported by the Institute of Mathematical Sciences of the Chinese University of Hong Kong and NSF grant 0196187.

The second author was supported by the NSA, NSF and the Clay Foundation.

In recent years it has become apparent that various arithmetic questions about higher dimensional algebraic varieties defined over number fields are also closely related to the cone of effective divisors. For example, given a variety X over a number field F, a line bundle L in the interior of NE^{1}(X), an open U ⊂ X over which L^{N}(N ≫ 0) is globally generated, and a height H_{L} associated to some adelic metrization L of L, we can consider the asymptotic behavior of the counting function

N(U, L, B) = #{x ∈ U(F) | H_{L}(x) ≤ B} B > 0.

There is a heuristic principle that, after suitably restricting U,

N(U, L, B) = c(L)B^{a(L)} log(B)^{b(L)−1}(1 + o(1)),

as B → ∞ (see [BT]). Here

a(L) := inf{a ∈ ℝ | aL + K_{X} ∈ NE^{1}(X)},

b(L) is the codimension of the face of NE^{1}(X) containing a(L)L+K_{X} (provided that NE^{1}(X) is locally polyhedral at this point), and c(L) > 0 is a constant depending on the chosen height (see [BM] and [BT] for more details). Notice that the explicit determination of the constant c(L) also involves the knowledge of the effective cone.

Such asymptotic formulas can be proved for smooth complete intersections in ℙ^{n} of small degree using the classical circle method in analytic number theory and for varieties closely related to linear algebraic groups, like flag varieties, toric varieties etc., using adelic harmonic analysis ([BT] and references therein). No general techniques to treat arbitrary varieties with many rational points are currently available. To our knowledge, the only other variety for which such an asymptotic is known to hold is the moduli space \overline{M}_{0,5} (Del Pezzo surface of degree 5) in its anticanonical embedding [dB]. Upper and lower bounds, with the expected a(L) and b(L), are known (see [VW]) for the Segre cubic threefold

Seg = {(x_{0}, …, x_{5}) : ∑_{j=0}^{5} x_{j}^{3} = ∑_{j=0}^{5} x_{j} = 0}.

This admits an explicit resolution by the moduli space \overline{M}_{0,6} (Remark 3.1); see [Hu] for the relationship between the Segre cubic and moduli spaces.

Our main result (Theorem 5.1) is a computation of the effective cone of \overline{M}_{0,6}. Besides the boundary divisors, the generators are the loci in \overline{M}_{0,6} fixed under

σ = (i_{1}i_{2})(i_{3}i_{4})(i_{5}i_{6}) ∈ 𝔖_{6}, {i_{1}, i_{2}, i_{3}, i_{4}, i_{5}, i_{6}} = {1, 2, 3, 4, 5, 6}.

This equals the closure of r^{∗}𝔥 ∩ M_{0,6}, where 𝔥 is the hyperelliptic locus in \overline{M}_{3}. The effective and moving cones of \overline{M}_{3} are studied in detail by Rulla [Ru]. Rulla’s inductive analysis of the moving cone is similar to the method outlined in Section 2. Results on the ample cone of \overline{M}_{0,6} have been recently obtained by Farkas and Gibney [FG].

The arithmetic consequences of Theorem 5.1 will be addressed in a future paper.

Acknowledgments: We are grateful to Seán Keel for helpful discussions, especially concerning the moving cone of \overline{M}_{0,6}, and to William Rulla, for pointing out an error in an early version of this paper. The discussion of the examples of ℚ-Fano fibrations benefitted from a remark by Gavril Farkas.

2. Generalities on effective cones

Let X be a nonsingular projective variety with Néron-Severi group NS(X) and group of one-cycles N_{1}(X). The closed effective cone of X is the closed convex cone

NE^{1}(X) ⊂ NS(X) ⊗ ℝ

generated by effective divisors on X. Let NM_{1}(X) be the dual cone NE^{1}(X)^{∗} in N_{1}(X) ⊗ ℝ. Similarly, let NE_{1}(X) be the cone of effective curves and NM^{1}(X) its dual, the nef cone.

We review one basic strategy, used in Section 5, for computing NE^{1}(X). Suppose we are given a collection Γ = {A_{1}, …, A_{m}} of effective divisors that we expect to generate the effective cone and a subset Σ ⊂ Γ. For any effective divisor E, we have a decomposition

E = M_{Σ} + B_{Σ}, B_{Σ} = a_{1}A_{1} + … + a_{m}A_{m}, a_{j} ≥ 0,

where B_{Σ} is the fixed part of |E| supported in Σ. The divisor M_{Σ} may have fixed components, but they are not contained in Σ. Let Mov(X)_{Σ} denote the closed cone generated by effective divisors without fixed components in Σ. To show that Γ generates NE^{1}(X) it suffices to show that it generates Mov(X)_{Σ}. Any divisor of Mov(X)_{Σ} restricts to an effective divisor on each A_{j} ∈ Σ. Consequently,

Mov(X)_{Σ} ⊂ NM_{1}(Σ, X)^{∗},

where NM_{1}(Σ, X) ⊂ N_{1}(X) is generated by the images of the NM_{1}(A_{i}) and A_{i} ∈ Σ. To prove that Γ generates NE^{1}(X), it suffices then to check that

{cone generated by Γ}^{∗} ⊂ NM_{1}(Σ, X).
3. Geometry of \overline{M}_{0,n}

3.1. A concrete description of \overline{M}_{0,n}. In this section we give a basis for the Néron-Severi group of \overline{M}_{0,n} and write down the boundary divisors and the symmetric group action.

We recall the explicit iterated blow-up realization

β_{n} : \overline{M}_{0,n} → ℙ^{n−3}

from [Has] (see also a related construction in [Kap].) This construction involves choosing one of the marked points; we choose s_{n}. Fix points p_{1}, …, p_{n−1} in linear general position in ℙ^{n−3} := X_{0}[n]. Let X_{1}[n] be the blow-up of ℙ^{n−3} at p_{1}, …, p_{n−1}, and let E_{1}, …, E_{n−1} denote the exceptional divisors (and their proper transforms in subsequent blow-ups). Consider the proper transforms ℓ_{ij} ⊂ X_{1}[n] of the lines joining p_{i} and p_{j}. Let X_{2}[n] be the blow-up of X_{1}[n] along the ℓ_{ij}, with exceptional divisors E_{ij}. In general, X_{k}[n] is obtained from X_{k−1}[n] by blowing-up along proper transforms of the linear spaces spanned by k-tuples of the points. The exceptional divisors are denoted

E_{i_{1},...,i_{k}} {i_{1}, …, i_{k}} ⊂ {1, …, n − 1}.

This process terminates with a nonsingular variety X_{n−4}[n] and a map

β_{n} : X_{n−4}[n] → ℙ^{n−3}.

One can prove that X_{n−4}[n] is isomorphic to \overline{M}_{0,n}. We remark that for a generic point p_{n} ∈ ℙ^{n−3}, we have an identification

β_{n}^{−1}(p_{n}) = (C, p_{1}, p_{2}, …, p_{n}),

where C is the unique rational normal curve of degree n − 3 containing p_{1}, …, p_{n} (see [Kap] for further information).

Let L be the pull-back of the hyperplane class on ℙ^{n−3} by β_{n}. We obtain the following explicit basis for NS(\overline{M}_{0,n}):

{L, E_{i_{1}}, E_{i_{1}i_{2}}, …, E_{i_{1},...,i_{k}}, …, E_{i_{1},...,i_{n−4}}}.

We shall use the following dual basis for the one-cycles N_{1}(\overline{M}_{0,n}):

{L^{n−4}, (−E_{i_{1}})^{n−4}, …, (−E_{i_{1},...,i_{k}})^{n−3−k}L^{k−1}, …, (−E_{i_{1},...,i_{n−4}})L^{n−5}}. (†)

3.2. Boundary divisors. Our next task is to identify the boundary divisors of \overline{M}_{0,n} in this basis. These are indexed by partitions

{1, 2, …, n} = S ∪ S^{c}, n ∈ S and |S|, |S^{c}| ≥ 2;

the generic point of the divisor D_{S} corresponds to a curve consisting of two copies of ℙ^{1} intersecting at a node ν, with marked points from S on one component and from S^{c} on the other. Thus we have an isomorphism

D_{S} ≃ \overline{M}_{0,|S|+1} × \overline{M}_{0,|S^{c}|+1},  (‡)
(ℙ^{1}, S) ∪_{ν} (ℙ^{1}, S^{c}) ⟶ (ℙ^{1}, S ∪ {ν}) × (ℙ^{1}, S^{c} ∪ {ν}).

The exceptional divisors are identified as follows:

E_{i_{1},...,i_{k}} = D_{i_{1},...,i_{k},n}, {i_{1}, …, i_{k}} ⊂ {1, …, n − 1}, k ≤ n − 4.

The remaining divisors D_{i_{1},...,i_{n−3},n} are the proper transforms of the hyperplanes spanned by (n − 3)-tuples of points; we have

[D_{i_{1},...,i_{n−3},n}] = L − E_{i_{1}} − E_{i_{2}} − … − E_{i_{1},...,i_{n−4}} − … − E_{i_{2},...,i_{n−3}}.

Remark 3.1. The explicit resolution of the Segre threefold

R : \overline{M}_{0,6} → Seg

alluded to in the introduction is given by the linear series

|2L − E_{1} − E_{2} − E_{3} − E_{4} − E_{5}|.

The image is a cubic threefold with ten ordinary double points, corresponding to the lines ℓ_{ij} contracted by R.

3.3. The symmetric group action on \overline{M}_{0,n}. The symmetric group 𝔖_{n} acts on \overline{M}_{0,n} by the rule

σ(C, s_{1}, …, s_{n}) = (C, s_{σ(1)}, …, s_{σ(n)}).

Let F_{σ} ⊂ \overline{M}_{0,n} denote the closure of the locus in M_{0,n} fixed by an element σ ∈ 𝔖_{n}.

We make explicit the 𝔖_{n}-action in terms of our blow-up realization. Choose coordinates (z_{0}, z_{1}, z_{2}, …, z_{n−3}) on ℙ^{n−3} so that

p_{1} = (1, 0, …, 0), …, p_{n−2} = (0, …, 0, 1), p_{n−1} = (1, 1, …, 1, 1).

Each permutation of the first (n − 1) points can be realized by a unique element of PGL_{n−2}. For elements of 𝔖_{n} fixing n, the action on \overline{M}_{0,n} is induced by the corresponding linear transformation on ℙ^{n−3}. Now let σ = (jn) and consider the commutative diagram

\begin{matrix}\overline{M}_{0,n} & →^{σ} & \overline{M}_{0,n} \\ β_{n} ↓ & & ↓ β_{n} \\ ℙ^{n−3} & −−→^{σ′} & ℙ^{n−3} \end{matrix} .

The birational map σ′ is the Cremona transformation based at the points p_{i_{1}}, …, p_{i_{n−2}} where

{i_{1}, …, i_{n−2}, j} = {1, 2, …, n − 1},

e.g., when σ = (n − 1, n) we have

σ(z_{0}, z_{1}, …, z_{n−3}) = (z_{1}z_{2} … z_{n−3}, z_{0}z_{2} … z_{n−3}, …, z_{0} … z_{n−4}).
4. Analysis of surfaces in \overline{M}_{0,6}
4.1. The \overline{M}_{0,5} case.

Proposition 4.1. NE^{1}(\overline{M}_{0,5}) is generated by the divisors D_{ij}, where \{ij\} ⊂ {1, 2, 3, 4, 5}. Sketch proof: This is well-known, but we sketch the basic ideas to introduce notation we will require later. As we saw in § 3.1, \overline{M}_{0,5} is the blow-up of ℙ^{2} at four points in general position. Consider the set of boundary divisors

Σ = {D_{i5}, D_{ij}} = {E_{i}, L − E_{i} − E_{j}}, {i, j} ⊂ {1, 2, 3, 4}

and the set of semiample divisors

Ξ = {L − E_{i}, 2L − E_{1} − E_{2} − E_{3} − E_{4}, L, 2L − E_{i} − E_{j} − E_{k}}, {i, j, k} ⊂ {1, 2, 3, 4}.

These semiample divisors come from the forgetting maps

ϕ_{i} : \overline{M}_{0,5} → \overline{M}_{0,4} ≃ ℙ^{1}, i = 1, …, 5

and the blow-downs

β_{i} : \overline{M}_{0,5} → ℙ^{2}, i = 1 …, 5.

Kleiman’s criterion yields

C(Σ) ⊂ NE_{1}(\overline{M}_{0,5}) = NM^{1}(\overline{M}_{0,5})^{∗} ⊂ C(Ξ)^{∗}.

All the inclusions are equalities because the cones generated by Ξ and Σ are dual; this can be verified by direct computation (e.g., using the computer program PORTA [PORTA]). □

4.2. Fixed points and the Cayley cubic. We identify the fixed-point divisors for the 𝔖_{6}-action on \overline{M}_{0,6}. When τ = (12)(34)(56) we have

τ(z_{0}, z_{1}, z_{2}, z_{3}) = (z_{0}z_{2}z_{3}, z_{1}z_{2}z_{3}, z_{0}z_{1}z_{2}, z_{0}z_{1}z_{3})

and F_{τ} is given by z_{0}z_{1} = z_{2}z_{3}. It follows that

[F_{τ}] = 2L − E_{1} − E_{2} − E_{3} − E_{4} − E_{5} − E_{13} − E_{23} − E_{24} − E_{14}.

More generally, when τ = (ab)(cd)(j6) we have

[F_{τ}] = 2L − E_{1} − E_{2} − E_{3} − E_{4} − E_{5} − E_{ac} − E_{ad} − E_{bc} − E_{bd}.

Remark 4.2. Consider (ℙ^{1}, s_{1}, …, s_{6}) ∈ F_{τ} and the quotient under the corresponding involution

q : ℙ^{1} ⟶ ℙ^{1}, q(s_{1}) = q(s_{2}), q(s_{3}) = q(s_{4}), etc.

Consider the map r : \overline{M}_{0,6} → \overline{M}_{3} identifying the pairs (12), (34), and (56) and write C = q(ℙ^{1}, s_{1}, …, s_{6}), so there is an induced q′ : C → ℙ^{1}. Thus C is hyperelliptic and F_{τ} corresponds to the closure of r^{∗}𝔥 ∩ \overline{M}_{0,6}, where 𝔥 ⊂ \overline{M}_{3} is the hyperelliptic locus.

Figure 1. Trinodal hyperelliptic curves

In Section 5.3 we will use the description of the effective cone of the fixed point divisors F_{σ}. We have seen that these are isomorphic to ℙ^{1} × ℙ^{1} blown-up at five points p_{1}, …, p_{5}. The projection from p_{5}

ℙ^{3} −−→ ℙ^{2}

induces a map φ : F_{σ} → ℙ^{2}, realizing F_{σ} as a blow-up of ℙ^{2}: Take four general lines ℓ_{1}, …, ℓ_{4} in ℙ^{2} with intersections q_{ij} = ℓ_{i} ∪ ℓ_{j}, and blow-up ℙ^{2} along the q_{ij}. We write

NS(F_{σ}) = ℤH + ℤG_{12} + … + ℤG_{34},

where the G_{ij} are the exceptional divisors and H is the pull back of the hyperplane class from ℙ^{2}.

Proposition 4.3. NE^{1}(F_{σ}) is generated by the (−1)-curves

G_{12}, …, G_{34}, H − G_{ij} − G_{kl},

and the (−2)-curves

H − G_{ij} − G_{ik} − G_{il}, {i, j, k, l} = {1, 2, 3, 4}.

Proof: Let Σ be the above collection of 13 curves. Consider also the following collection Ξ of 38 divisors, grouped as orbits under the 𝔖_{4}-action:

typical member orbit size

induced morphism

H  1  blow-down φ : F_{σ} → ℙ^{2}
H − G_{12}  6  conic bundle F_{σ} → ℙ^{1}
2H − G_{12} − G_{13} − G_{23}  4  blow-down F_{σ} → ℙ^{2}
2H − G_{12} − G_{23} − G_{34}  12  blow-down F_{σ} → ℙ^{2}
2H − G_{12} − G_{23} − G_{34} − G_{14}  3  conic bundle F_{σ} → ℙ^{1}
3H − 2G_{12} − G_{13} − G_{23} − G_{34}  12  blow-down F_{σ} → ℙ(1, 1, 2)

Note that each of these divisors is semiample: the corresponding morphism is indicated in the table. In particular,

C(Σ) := {cone generated by Σ} ⊂ NE_{1}(F_{σ}),
C(Ξ) := {cone generated by Ξ} ⊂ NM^{1}(F_{σ})

and Kleiman’s criterion yields

C(Σ) ⊂ NE_{1}(F_{σ}) = NM^{1}(F_{σ})^{∗} ⊂ C(Ξ)^{∗}.

A direct verification using PORTA [PORTA] shows that the cones C(Σ) and C(Ξ) are dual, so all the inclusions are equalities. □

Remark 4.4. The image of F_{τ} under the resolution R of 3.1 is a cubic surface with four double points, classically called the Cayley cubic [Hu].

5. The effective cone of \overline{M}_{0,6}

We now state the main theorem:

Theorem 5.1. The cone of effective divisors NE^{1}(\overline{M}_{0,6}) is generated by the boundary divisors and the fixed-point divisors F_{σ}, where σ ∈ 𝔖_{6} is a product of three disjoint transpositions.

5.1. Proof of Main Theorem. We use the strategy outlined in § 2. Consider the collection of boundary and fixed-point loci

Γ = {D_{ij}, D_{ijk}, F_{σ}, σ = (ij)(kl)(ab), {i, j, k, l, a, b} = {1, 2, 3, 4, 5, 6}}

and the subset of boundary divisors

Σ = {D_{ij}, D_{ijk}}.

We compute the cone NM_{1}(Σ, \overline{M}_{0,6}), the convex hull of the union of the images of NM_{1}(D_{ij}) and NM_{1}(D_{ijk}) in N_{1}(\overline{M}_{0,6}). Throughout, we use the dual basis for N_{1}(\overline{M}_{0,6}) (cf. (†)):

{L^{2}, E_{1}^{2}, E_{2}^{2}, E_{3}^{2}, E_{4}^{2}, E_{5}^{2}, −LE_{12}, −LE_{13},
−LE_{14}, −LE_{15}, −LE_{23}, −LE_{24}, −LE_{25}, −LE_{34}, −LE_{35}, −LE_{45}}.

Recall the isomorphism (‡)

(π_{ijk}, π_{lab}) : D_{ijk} ⟶ ℙ^{1} × ℙ^{1}, {i, j, k, l, a, b, c} = {1, 2, 3, 4, 5, 6}

so that

N_{1}(D_{ijk}) = ℤB_{ijk} ⊕ ℤB_{lab}, NM_{1}(D_{ijk}) = ℝ_{+}B_{ijk} + ℝ_{+}B_{lab},

where B_{ijk} is the class of the fiber of π_{ijk} (and its image in N_{1}(\overline{M}_{0,6})). For example, the inclusion j_{345} : D_{345} ↪ \overline{M}_{6} induces

(j_{345})_{∗} = ( \begin{matrix}0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ 1 & 1 & 1 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \end{matrix} )^{T}

using the bases (†) for N_{1}(\overline{M}_{0,6}) and \{B_{126}, B_{345}\} for N_{1}(D_{345}). In particular, we find

NM_{1}(\{D_{ijk}\}, \overline{M}_{0,6}) = C(\{B_{ijk}\}), \{i, j, k\} ⊂ \{1, 2, 3, 4, 5, 6\},

with \binom{6}{3} = 20 generators permuted transitively by 𝔖_{6} (Table 1).

The boundary divisor D_{ij} is isomorphic to \overline{M}_{0,5} with marked points \{k, l, a, b, ν\} where {i, j, k, l, a, b} = {1, 2, 3, 4, 5, 6} and ν is the node (cf. formula (‡)). The proof of Proposition 4.1 gives generators for the nef cone of D_{ij}. Thus the cone NM_{1}(D_{ij}, \overline{M}_{0,6}) is generated by the classes

\{A_{ij}, A_{ij;k}, A_{ij;l}, A_{ij;a}, A_{ij;b}, C_{ij}, C_{ij;k}, C_{ij;l}, C_{ij;a}, C_{ij;b}\} ⊂ N_{1}(\overline{M}_{0,6})

corresponding to the forgetting and blow-down morphisms

{ϕ_{ν}, ϕ_{k}, ϕ_{l}, ϕ_{a}, ϕ_{b}, β_{ν}, β_{k}, β_{l}, β_{a}, β_{b}}.

As an example, consider the inclusion j_{45} : D_{45} ↪ \overline{M}_{0,6} with

j_{45}_{∗} = ( \begin{matrix}1 & 0 & 0 & 0 & 0 & 0 & 1 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 \\ 0 & 1 & 0 & 0 & 0 & 0 & −1 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ 0 & 0 & 1 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 \\ 0 & 0 & 0 & 1 & 0 & 0 & 0 & −1 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 \\ 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 \end{matrix} )^{T} .

Applying this to the nef divisors of D_{45} gives the generators for NM_{1}(D_{45}, \overline{M}_{0,6}) (Table 2).

However, four of the (−1)-curves in D_{ij} are contained in D_{ijk}, D_{ijl}, D_{ija}, and D_{ijb}, with classes B_{ijk}, B_{ijl}, B_{ija}, and B_{ijb} respectively. Thus we have the relations

C_{ij} = A_{ij;k} + B_{ijk}, C_{ij;k} = A_{ij} + B_{ijk}

which implies that the C_{ij} and C_{ij;k} are redundant:

Proposition 5.2. The cone NM_{1}(Σ, \overline{M}_{0,6}) is generated by the A_{ij}, the A_{ij;k}, and the B_{ijk}.

These are written out in Tables 1,3, and 4.

Our next task is to write out the generators for the dual cone C(Γ)^{∗}, as computed by PORTA [PORTA]. Since Γ is stable under the 𝔖_{6} action, so are C(Γ) and its dual cone. For the sake of brevity, we only write 𝔖_{6}-representatives of the generators, ordered by anticanonical degree.

The discussion of Section 2 shows that Theorem 5.1 will follow from the inclusion

C(Γ)^{∗} ⊂ NM_{1}(Σ, \overline{M}_{0,6}).

We express each generator of C(Γ)^{∗} as a sum (with non-negative coefficients) of the {A_{ij}, A_{ij;k}, B_{ijk}}. Both cones are stable under the 𝔖_{6}-action, so it suffices to produce expressions for one representative of each 𝔖_{6}-orbit. We use the representatives from Table 5:

(1) = A_{15} + A_{13} + A_{35} + 2B_{246} (2) = A_{34;5} + B_{126}
(3) = A_{15} + A_{14} + 2B_{236} (4) = A_{25} + B_{146} + B_{136}
(5) = A_{23} + B_{146} + B_{156} + B_{236} (6) = A_{15} + A_{14} + B_{236} + B_{246} + B_{356}
(7) = A_{13;5} + A_{15} + B_{236} + B_{246} (8) = A_{12;5} + A_{14} + B_{256} + B_{356}
(9) = A_{24} + A_{34} + B_{126} + B_{136} + B_{156}
(10) = A_{25} + B_{136} + B_{146} + B_{256} + B_{346}
(11) = A_{34;5} + A_{35;4} + A_{25;3} + B_{146} (12) = A_{12} + A_{34} + 2B_{126} + 2B_{346}
(13) = A_{15} + A_{14} + A_{23} + 2B_{146} + 2B_{236}
(14) = A_{23;5} + A_{15} + A_{25} + B_{136} + B_{146} + B_{236}
(15) = A_{23;5} + A_{24;5} + A_{15} + B_{156} + B_{346}
(16) = A_{24;5} + A_{15} + B_{136} + B_{156} + B_{236}
(17) = A_{23} + 2A_{25} + 2B_{136} + 2B_{146}
(18) = A_{12;3} + A_{34} + B_{126} + B_{136} + A_{36;1}
(19) = A_{12;5} + A_{15} + A_{25} + B_{136} + B_{246} + B_{346}
(20) = A_{13;5} + A_{35} + A_{45} + 2B_{126} + B_{456}
(21) = A_{12} + A_{13} + B_{126} + B_{136} + B_{246} + B_{346} + B_{456}
(22) = A_{15} + A_{23} + A_{34} + B_{126} + B_{146} + B_{156} + 2B_{236}
(23) = A_{13;5} + A_{14;5} + A_{23} + A_{13} + B_{256} + 2B_{456}
(24) = A_{15} + A_{23} + A_{24} + A_{34} + B_{126} + B_{136} + B_{146} + B_{156} + 2B_{236}
(25) = 2A_{14} + A_{24} + 2A_{13;5} + 2B_{256} + 2B_{356}

This completes the proof of Theorem 5.1.□

5.2. Geometric interpretations of coextremal rays. By definition, a coextremal ray ℝ_{+}ρ ⊂ NM_{1}(X) satisfies the following

• for any nontrivial ρ_{1}, ρ_{2} ∈ NM_{1}(X) with ρ_{1} + ρ_{2} ∈ ℝ_{+}ρ, ρ_{1}, ρ_{2} ∈ ℝ_{+}ρ;

• K_{X}ρ < 0.

Batyrev ([Ba], Theorem 3.3) shows that, for smooth (or ℚ-factorial terminal) threefolds, the minimal model program yields a geometric interpretation of coextremal rays. They arise from diagrams

\begin{aligned}X & −−→^{ψ} & Y \\ & & ↓ μ \\ & & B \end{aligned}

where ψ is a sequence of birational contractions and μ is a Mori fiber space. The coextremal ray ρ = ψ^{∗}[C], where C is a curve lying in the general fiber of μ. These interpretations will hold for higher-dimensional varieties, provided the standard conjectures of the minimal model program are true.

It is natural then to write down these Mori fiber space structures explicitly. Our analysis makes reference to the list of orbits of coextremal rays in Table 5 (and uses the same numbering):

(1) The first orbit in the Table is orthogonal to each of the boundary divisors D_{ij} ⊂ \overline{M}_{0,6}. The ℚ-Fano fibration associated with this coextremal ray must contract these divisors. The anticanonical series | − K_{\overline{M}_{0,6}}| yields a birational morphism

\overline{M}_{0,6} → J ⊂ ℙ^{4}

onto a quartic ℚ-Fano hypersurface, called the Igusa quartic [Hu]. The fifteen singular points of the Igusa quartic are the images of the D_{ij}. The coextremal ray has anticanonical degree two and corresponds to curves passing through the generic point, i.e., the conics in J.

(2) Forgetting any of the six marked points

\overline{M}_{0,6} → \overline{M}_{0,5}

yields a Mori fiber space, and the fibers are coextremal.

(3) We define a conic bundle structure on \overline{M}_{0,6} by explicit linear series, using the blow-up description of Subsection 3.1. Consider the cubic surfaces in ℙ^{3} passing through the lines

ℓ_{14}, ℓ_{15}, ℓ_{24}, ℓ_{25}, ℓ_{34}, ℓ_{35}.

This linear series has additional base points: Any cubic surface containing the lines ℓ_{14}, ℓ_{24}, ℓ_{34} (resp. ℓ_{15}, ℓ_{25}, ℓ_{35}) must be singular at p_{4} (resp. p_{5}), and thus contains the line ℓ_{45} by the Bezout Theorem.

Our linear series has projective dimension two. Indeed, cubic hypersurfaces in ℙ^{3} depend on 19 parameters; the singularities at p_{4} and p_{5} each impose four conditions, the remaining points p_{1}, p_{2}, p_{3} impose three further conditions, and containing the six lines imposes six more conditions. Thus we obtain a conic bundle structure

μ : \overline{M}_{0,6} −−→ ℙ^{2}

collapsing the two-parameter family of conics passing through the six lines above.

(4) For any two disjoint subsets {i, j}, {k, l} ⊂ {1, 2, 3, 4, 5, 6} we consider the forgetting maps

ϕ_{ij} : \overline{M}_{0,6} → ℙ^{1}, ϕ_{kl} : \overline{M}_{0,6} → ℙ^{1}.

Together, these induce a conic bundle structure

(ϕ_{ij}, ϕ_{kl}) : \overline{M}_{0,6} → ℙ^{1} × ℙ^{1}.

The class of a generic fiber is coextremal.

5.3. The moving cone. Our analysis gives, implicitly, the moving cone of \overline{M}_{0,6}:

Theorem 5.3. The closed moving cone of \overline{M}_{0,6} is equal to NM_{1}(Γ, \overline{M}_{0,6})^{∗}, where Γ is the set of generators for NE^{1}(\overline{M}_{0,6}).

In the terminology of [Ru], the ‘inductive moving cone’ equals the ‘moving cone’. Combining Theorem 5.3 with the computation of the ample cones to the boundaries D_{ij} and D_{ijk} and the fixed-point divisors F_{σ} (Proposition 4.3) we obtain the moving cone. However, finding explicit generators for the moving cone is a formidable computational problem.

Proof: Recall that \overline{M}_{0,6} is a log Fano threefold: −(K_{\overline{M}_{0,6}} + ϵ ∑_{ij} D_{ij}) is ample for small ϵ > 0 [KeMc]. Using Corollary 2.16 of [KeHu], it follows that \overline{M}_{0,6} is a ‘Mori Dream Space’. The argument of Theorem 3.4.4 of [Ru] shows that an effective divisor on \overline{M}_{0,6} that restricts to an effective divisor on each generator A_{i} ∈ Γ is in the moving cone. □

Remark 5.4. Our proof of Theorem 5.1 uses the cone NM_{1}(Σ, \overline{M}_{0,6})^{∗}, rather than the (strictly) smaller moving cone. Of course, if the coextremal rays are in NM_{1}(Σ, \overline{M}_{0,6}), a fortiori they are in NM_{1}(Γ, \overline{M}_{0,6}).

References

[Ba] V.V. Batyrev, The cone of effective divisors of threefolds, Proceedings of the International Conference on Algebra, Part 3 (Novosibirsk, 1989), 337–352, Contemp. Math., 131, Part 3, Amer. Math. Soc., Providence, RI, 1992.

[BM] V. V. Batyrev and Yu. I. Manin, Sur le nombre des points rationnels de hauteur borné des variétés algébriques, Math. Ann. 286 (1990), no. 1-3, 27–43.

[BT] V. V. Batyrev and Y. Tschinkel, Tamagawa numbers of polarized algebraic varieties, Nombre et répartition de points de hauteur bornée (Paris, 1996), Astérisque 251 (1998), 299–340.

[dB] R. de la Bretèche, Nombre de points de hauteur bornée sur les surfaces de del Pezzo de degré 5, preprint, Université Paris-Sud (2000).

[FG] G. Farkas and A. Gibney, The nef cone of moduli spaces of pointed curves of small genus, in preparation.

[F] T. Fujita, On Kodaira energy and adjoint reduction of polarized manifolds, Manuscripta Math. 76 (1992), no. 1, 59–84.

[Has] B. Hassett, Moduli spaces of weighted pointed stable curves, preprint (2001).

[Hu] B. Hunt, The geometry of some special arithmetic quotients, Lecture Notes in Mathematics 1637, Springer-Verlag, Berlin, 1996.

[Kap] M. M. Kapranov, Veronese curves and Grothendieck-Knudsen moduli space \overline{M}_{0,n}, J. Algebraic Geom. 2 (1993), no. 2, 239–262.

[Ke] S. Keel, personal communication, 1999.

[KeHu] Y. Hu and S. Keel, Mori dream spaces and GIT, Michigan Math. J. 48 (2000), 331–348.

[KeMc] S. Keel and J. McKernan, Contraction of extremal rays on \overline{M}_{0,n}, alg-geom 9607009.

[PORTA] Polyhedron  representation  transformation  algorithm,  available  at

 http://www.iwr.uni-heidelberg.de/groups/comopt/software/PORTA.

[Ru] W. Rulla, The Birational Geometry of \overline{M}_{3} and \overline{M}_{2,1}, Ph.D. Thesis, University of Texas at Austin (2001).

[VW] R. C. Vaughan and T. D. Wooley, On a certain nonary cubic form and related equations, Duke Math. J. 80 (1995), no. 3, 669–735.

[Ve] P. J. Vermeire, A Counter Example to Fulton’s Conjecture on \overline{M}_{0,n}, to appear in the Journal of Algebra.

Department of Mathematics, Rice University, Houston, Texas 77251

    E-mail address: hassett@math.rice.edu

Department of Mathematics, Princeton University, Princeton, NJ 08544-1000

    E-mail address: ytschink@math.princeton.edu

Table 1. Generators for NM_{1}(\{D_{ijk}\}, \overline{M}_{0,6})

\begin{matrix}B_{126} & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ B_{136} & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ B_{146} & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ B_{156} & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 \\ B_{236} & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 \\ B_{246} & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 \\ B_{256} & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 \\ B_{346} & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 \\ B_{356} & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 \\ B_{456} & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 \\ B_{123} & 1 & 0 & 0 & 0 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 \\ B_{124} & 1 & 0 & 0 & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 \\ B_{125} & 1 & 0 & 0 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 \\ B_{134} & 1 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 \\ B_{135} & 1 & 0 & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 \\ B_{145} & 1 & 0 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 \\ B_{234} & 1 & 1 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 \\ B_{235} & 1 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ B_{245} & 1 & 1 & 0 & 1 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ B_{345} & 1 & 1 & 1 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \end{matrix}

Table 2. Generators for NM_{1}(D_{45}, \overline{M}_{0,6})

\begin{matrix}A_{45} & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 1 \\ A_{45;1} & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 \\ A_{45;2} & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{45;3} & 1 & 0 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{45;6} & 2 & 1 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 \\ C_{45;6} & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 \\ C_{45;1} & 2 & 0 & 1 & 1 & 0 & 0 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 \\ C_{45;2} & 2 & 1 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 1 \\ C_{45;3} & 2 & 1 & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 1 \\ C_{45} & 2 & 1 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \end{matrix}

Table 3. Generators A_{ij} for NM_{1}(\{D_{ij}\}, \overline{M}_{0,6})

\begin{matrix}A_{12} & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 1 & 1 \\ A_{13} & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 1 & 1 & 0 & 0 & 1 \\ A_{14} & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 1 & 0 \\ A_{15} & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 1 & 1 & 0 & 1 & 0 & 0 \\ A_{16} & 0 & −2 & 0 & 0 & 0 & 0 & 1 & 1 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{23} & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 1 & 1 & 0 & 0 & 0 & 0 & 1 \\ A_{24} & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 1 & 0 & 1 & 0 & 0 & 1 & 0 \\ A_{25} & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 1 & 0 & 0 & 0 & 1 & 1 & 0 & 0 \\ A_{26} & 0 & 0 & −2 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 1 & 1 & 1 & 0 & 0 & 0 \\ A_{34} & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 1 & 1 & 0 & 0 \\ A_{35} & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 1 & 0 \\ A_{36} & 0 & 0 & 0 & −2 & 0 & 0 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 1 & 1 & 0 \\ A_{45} & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 1 \\ A_{46} & 0 & 0 & 0 & 0 & −2 & 0 & 0 & 0 & 1 & 0 & 0 & 1 & 0 & 1 & 0 & 1 \\ A_{56} & 0 & 0 & 0 & 0 & 0 & −2 & 0 & 0 & 0 & 1 & 0 & 0 & 1 & 0 & 1 & 1 \end{matrix}

Table 4. Generators A_{ij;k} for NM_{1}(\{D_{ij}\}, \overline{M}_{0,6})

\begin{matrix}A_{12;3} & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 \\ A_{12;4} & 1 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 \\ A_{12;5} & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 \\ A_{12;6} & 2 & 0 & 0 & 1 & 1 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{13;2} & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 \\ A_{13;4} & 1 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 \\ A_{13;5} & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 \\ A_{13;6} & 2 & 0 & 1 & 0 & 1 & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{14;2} & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 \\ A_{14;3} & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 \\ A_{14;5} & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 \\ A_{14;6} & 2 & 0 & 1 & 1 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{15;2} & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 \\ A_{15;3} & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 \\ A_{15;4} & 1 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 \\ A_{15;6} & 2 & 0 & 1 & 1 & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{16;2} & 0 & −1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{16;3} & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{16;4} & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{16;5} & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{23;1} & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 \\ A_{23;4} & 1 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{23;5} & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{23;6} & 2 & 1 & 0 & 0 & 1 & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 \\ A_{24;1} & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 \\ A_{24;3} & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{24;5} & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{24;6} & 2 & 1 & 0 & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 \\ A_{25;1} & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 \\ A_{25;3} & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{25;4} & 1 & 0 & 0 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{25;6} & 2 & 1 & 0 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 \\ A_{26;1} & 0 & 0 & −1 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{26;3} & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 \\ A_{26;4} & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 \\ A_{26;5} & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 \\ A_{34;1} & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 \\ A_{34;2} & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{34;5} & 1 & 0 & 0 & 0 & 0 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{34;6} & 2 & 1 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 \\ A_{35;1} & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 \\ A_{35;2} & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{35;4} & 1 & 0 & 0 & 0 & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{35;6} & 2 & 1 & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 \\ A_{36;1} & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{36;2} & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 \\ A_{36;4} & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 \\ A_{36;5} & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 \\ A_{45;1} & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 \\ A_{45;2} & 1 & 0 & 1 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{45;3} & 1 & 0 & 0 & 1 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{45;6} & 2 & 1 & 1 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 \\ A_{46;1} & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{46;2} & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 \\ A_{46;3} & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 \\ A_{46;5} & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 \\ A_{56;1} & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 1 & 0 & 0 & 0 & 0 & 0 & 0 \\ A_{56;2} & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 & 0 & 0 \\ A_{56;3} & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 & 0 \\ A_{56;4} & 0 & 0 & 0 & 0 & 0 & −1 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 0 & 1 \end{matrix}

Table 5. 𝔖_{6}-orbits of coextremal rays of \overline{M}_{0,6}

deg_{−K} order
(1)  2  1 3 0 0 0 0 0 1 1 1 1 1 1 1 1 1 1
(2)  2  6 1 0 0 0 0 1 0 0 0 0 0 0 0 0 0 0
(3)  2  15 2 0 0 0 0 0 0 0 1 1 0 1 1 1 1 0
(4)  2  45 1 0 0 0 0 0 0 0 0 0 0 0 1 1 0 0
(5)  3  60 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1
(6)  3  72 2 0 0 0 0 0 0 0 1 1 1 0 1 1 0 0
(7)  3  120 2 0 0 0 0 1 0 0 0 1 0 1 0 1 0 0
(8)  3  120 2 0 0 0 0 1 0 0 1 0 1 0 0 1 0 0
(9)  3  180 2 0 0 0 0 0 0 0 0 1 0 1 1 1 1 0
(10)  4  6 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
(11)  4  10 3 0 0 1 1 1 2 0 0 0 0 0 0 0 0 0
(12)  4  30 2 0 0 0 0 0 0 0 0 1 0 0 1 0 1 1
(13)  4  60 3 0 0 0 0 0 0 0 0 2 1 1 1 1 1 1
(14)  4  90 3 0 0 0 0 1 0 0 1 1 0 1 1 2 0 0
(15)  4  90 3 0 0 0 0 2 0 1 1 0 1 1 0 0 0 0
(16)  4  180 2 0 0 0 0 1 0 0 0 0 0 1 0 1 0 0
(17)  4  180 3 0 0 0 0 0 0 0 1 1 1 0 2 2 0 1
(18)  4  360 2 0 0 0 0 0 0 0 0 1 0 0 1 1 0 1
(19)  4  360 3 0 0 0 0 1 0 0 1 1 1 0 1 2 0 0
(20)  4  360 3 0 0 0 0 1 0 1 1 0 1 2 0 0 1 0
(21)  5  120 2 0 0 0 0 0 0 0 0 0 0 0 1 0 1 1
(22)  5  360 3 0 0 0 0 0 0 0 0 2 0 1 1 2 0 1
(23)  5  360 4 0 0 0 0 2 0 1 1 1 2 2 0 0 0 0
(24)  6  360 4 0 0 0 0 0 0 0 0 3 0 2 1 2 1 1
(25)  6  360 5 0 0 0 0 2 0 1 2 1 2 3 0 0 1 0